OpenStudy (anonymous):

Simplify the expression log(12)...base is 10...how do you find without using a calculator?

OpenStudy (anonymous):

Do they tell us any decimal values for some other logs? Sometimes problems also give that.

OpenStudy (debbieg):

Are you sure they want it "simplified"? or maybe written as a combination other logs, e.g, log(2) and log(3)?

OpenStudy (anonymous):

If now, we can do some things to simplify it. The 12 is not any power of 10, so it's not going to give us a nice whole number answer or a fraction. But we can do is use some properties of logarithms to break it down.

OpenStudy (anonymous):

if not*

OpenStudy (wolf1728):

log (square root of 10) =.5 That has nothing to do with the problem but at least it gives you some idea as to which numbers convert to logs easily.

OpenStudy (anonymous):

maybe i wrote it wrong

OpenStudy (anonymous):

One of them tells us that a multiplication becomes an addition of logarithms. Log xy= log x + log y Have you see that before?

OpenStudy (anonymous):

Yes, probably. Please check!

OpenStudy (wolf1728):

Thanks Debbie - it's been corrected

OpenStudy (anonymous):

using the properties of logarithms with numerical values. Find the logarithm

OpenStudy (anonymous):

the examples they give are easy...then they throw a boomerang and me

OpenStudy (anonymous):

and = at

OpenStudy (anonymous):

oh ok

OpenStudy (debbieg):

Here are properties that you might use for this problem: \[\Large \log(MN)=\log M+\log N\] \[\Large \log M^a=a\log M\]

OpenStudy (anonymous):

It says numerically values, so I'm thinking somewhere they tell us something like log3=... or log 2 = ...

OpenStudy (anonymous):

\[\log_{10} 12\] without using a calculator

OpenStudy (debbieg):

Can you tell us one of the examples?

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

Here's all I think we can really do. 12 breaks down into 3 times 4

OpenStudy (anonymous):

log5(25) i know this is 2 which is not that hard...nothing complicated

OpenStudy (anonymous):